One card is selected at random from a deck of cards. Determine the probability that the card is a red 10.
Answer by stanbon(75887)

One card is selected at random from a deck of cards. Determine the probability that the card is a red 10.
---
There are 2 red 10's
----
P(select a red 10) = 2/52 = 1/26
=================================
Cheers,
Stan H.